Two identical cells each of emf E and internal resistance r are connected in parallel with an external resistance R. To get maximum
power developed across R, the value of R is
(A) R = r/2 (B) R = r (C) R = r/3 (D) R = 2r

Grade:

2 Answers

Arun
3 years ago
Equivalent resistance Req=2r+R=2r+2R
I=r+2R2E
 
For maximum power consumption I should be maximum so denominator is minimum. For this
r+2R=(r2R)2+2r2R
 
r2R=0
 
R=2r
Vikas TU
3 years ago
first of all You should know that parallel cells just increase the capacity of the cell(means life) but keeps the voltage same as of the single cell emf.
So for max. Power
P=E^2/4R (r=R)
Now for your question 2 cells have internal resistance of r and connected in parallel so equivalent internal resistance is r/2
So here is our answer
R=r/2.